Exploring the World Wide Web

John Kester and Marion French

Wallace Memorial Library, RIT

Abstract

The presentation will examine the multitude of Computer Science resources available through the WWW. We will use Netscape as our browsing tool to retrieve documents that contain text, graphics, and sound as well as links to other documents.

We will also search the resources available, both on-campus and offcampus, through the library’s InfoNet menu.

Note: This event will be held in ICL1, not in the conference room